### v1

@@ 2. Related Work
Despite these applications, to the best of our knowledge, the branching random walk remains underexplored in distributed computing and this work seeks to showcase its untapped potential.
3. Algorithm Overview
In this section, we provide a high-level overview of our algorithm, starting with its main building block: COBRA.
3.1. Coalescing-Branching Random Walk
@@
The COalescing-BRAnching Random Walk (COBRA walk) was first introduced by Dutta et al. (dutta2015coalescing, ) in their work ”Coalescing-Branching Random Walks on Graphs” (dutta2015coalescing, ), with subsequent refinements presented in (cooper2016coalescing, ; mitzenmacher2018better, ; berenbrin2018tight, ). The COBRA walk is a generalization of the classical random walk, defined as follows: At round 00, a source node $s\in V$ possesses a token. At round $r$, each node possessing a token selects $\kappa$ of its neighbors uniformly at random, sends a token copy to each of them, and these neighbors are said to possess a token at round $r+1$. Here, $\kappa$, referred to as the branching factor, can be generalized to any positive real number (cooper2016coalescing, ). When $\kappa=1$, the COBRA walk reduces to the classical random walk. From now on, we consider $\kappa$ to always be $2$. It is important to note that if a node receives multiple tokens in a round, it behaves as if it has received only one token; it will still choose $\kappa$ neighbors uniformly at random. This property, where received tokens coalesce at a node, gives the primitive its name.
@@
Cooper et al. (cooper2016coalescing, ) studied the cover time $T$ of the COBRA walk on regular expanders and obtained the following theorem which we use in our result
Theorem 1 (Cooper et al. (cooper2016coalescing, )).
@@
Let $G$ be a connected $n$-vertex regular graph. Let $\lambda_{2}$ be the second largest eigenvalue (in the absolute value) of the normalized adjacency matrix of $G$. Then
@@
with probability at least $1-O(\frac{1}{n^{2}})$.
Remark 2.
In the paper, the bound on probability is $1-O(\frac{1}{n})$, though the analysis, which is based on Chernoff’s bounds, can be adapted so that the probability is $1-O(n^{C})$ for any constant $C$ and the cover time is only multiplied by a constant.
In the upcoming analysis of our result, we will make sure that $\lambda_{2}$ is no more than $\frac{13}{14}$ w.h.p. Let $C_{T}$ be a sufficiently large constant so that a COBRA walk covers a regular graph with $\lambda_{2}\leq\frac{13}{14}$ with probability at least $1-O(\frac{1}{n^{2}})$ in $C_{T}\log n$ rounds. From now on, we define $T$ to be $C_{T}\log n$.
3.2. Algorithm Description
The algorithm proceeds through the following steps:
  1. (1)
    Building a BFS Tree and Gathering Information: Nodes construct a BFS tree rooted at source node $s$. Using the tree, every node learns the total number of nodes $|V|$, the minimum degree $\delta$, and the maximum degree $\Delta$. This step takes $O(D)$ rounds and does not require any prior knowledge of the graph topology.
  2. (2)
    Regularizing the Graph: Each node $v$ adds $\Delta-\deg(v)$ self-loops to its adjacency list to make the graph regular. We will show in our analysis that each node adds a relatively small number of self-loops. This operation is purely local and requires no communication between nodes.
  3. (3)
    Constructing Spanning Subgraphs through Multiple COBRA Walks: The source node $s$ initiates $\delta$ COBRA walks by creating $\delta$ tokens labeled $1$ through $\delta$. When a token from the $i$-th COBRA walk is sent along an edge $e:\ (u,v)$, $u$ and $v$ mark $e$ as part of the $i$-th subgraph. Note that a single edge may belong to multiple subgraphs.
    When running multiple COBRA walks simultaneously, congestion can occur if a node attempts to send multiple tokens from different COBRA walks along the same edge in a single round. Since only one token can traverse an edge per round, this creates a bottleneck that needs to be managed. To address this, we organize the process into phases, where each phase consists of $4\sqrt{\log n}$ rounds. During each phase, nodes handle the token distribution for a single round of all COBRA walks. We will show that the randomness in the process ensures that no edge is assigned more than $4\sqrt{\log n}$ tokens in a single phase, allowing the walks to proceed without interfering with each other.
@@ 3.2. Algorithm Description
  4. (4)
    Constructing Tree Packings: After $T$ phases, nodes stop sending tokens. The source $s$ initiates a BFS on each subgraph to transform it into a tree. By the end of this step, the algorithm constructs a tree packing $\{T_{i}\}_{i\in[\delta]}$. Since some edges might belong to multiple trees, this step takes the number of rounds that is at most the height of the highest tree times the maximal number of subgraphs a single edge belongs to, that is at most $T\cdot T\cdot 4\sqrt{\log n}=\tilde{O}(1)$.
  5. (5)
    Distributing Messages: The source node $s$ evenly divides the set of messages $M$ across the $\delta$ trees, ensuring that each tree receives $\frac{k}{\delta}$ messages. These messages are then downcasted along the trees one by one. To broadcast $k$ messages using the tree packing of size $\delta$, height $O(T)$ and weight $O(T\cdot\sqrt{\log n})$ we spend $O(T\cdot\sqrt{\log n}\cdot(\frac{k}{\delta}+T))$ rounds.
4. Proof
In this section, we formally state all our main results and their auxiliaries. We start by providing a high level proof.
4.1. Proof Outline
The proof proceeds in three main steps. First, we show that after making a random graph regular by adding self-loops, it retains its expansion properties. To achieve this, we use the following results. Theorem 3 by Hoffman et al. demonstrates that a random graph is a good expander w.h.p. Lemma 4, using Chernoff bounds, establishes that a random graph is almost regular w.h.p., that is $\frac{\Delta}{\delta}$ is close to $1$. Finally, Lemma 5, based on Weyl’s theorem, shows that small perturbations to the diagonal elements of a matrix result in small perturbations to its eigenvalues. Combining these results, we conclude that regularizing the graph does not significantly impact its expansion properties.
The second step is to prove that each individual COBRA walk successfully covers the entire network. Lemma 9 ensures that w.h.p., in each phase, every node successfully distributes its tokens, ensuring that no edge is overloaded with more than $4\sqrt{\log n}$ tokens. Using this, Theorem 1 by Cooper et al., together with a union bound, guarantees that all COBRA walks complete successfully within $O(\log n)$ phases w.h.p.
In the final step, we argue that the algorithm produces a tree packing with size $\delta(G)$, height $O(\log n)$, and weight $O(\log^{2.5}n)$. This tree packing allows for broadcasting all messages in $\tilde{O}(\frac{k}{\delta(G)})$ rounds. This is optimal up to a polylogarithmic factor, as $\frac{k}{\delta(G)}$ provides a natural lower bound for the broadcast time.
4.2. Introducing Regularity while Maintaining Expansion
@@
Theorem 3 (Hoffman et al. (hoffman2021spectral, )).
For a positive constant $C$ and $p\geq\frac{C\log n}{n}$, consider an Erdős–Rényi graph $G(n,p)$. Then, with probability at least $1-O(\frac{1}{n^{C-1}})$ we have $\lambda_{2}(G)=O(\frac{1}{\sqrt{pn}})$.
The following Lemma shows that with high probability an Erdős–Rényi graph $G(n,p)$ is almost regular.
Lemma 4.
@@
Proof.
Let us fix a vertex $v$ and consider the number of edges it might have. For each potential edge $e_{i}$, $i\in[n-1]$ let us introduce an indicator variable $\chi_{i}$ which is equal to $1$ if the edge exists and to 00 if it does not. The number of edges $v$ has is then $\sum\limits_{i\in[n-1]}\chi_{i}$. The expectation of that is $p(n-1)$ and applying Chernoff’s bounds we get
$Pr\left[deg(v)\geq(1+\frac{1}{15})C_{p}\log n\right]\leq\exp\left(-\frac{C_{p}\log n}{675}\right)\leq\frac{1}{2n^{3}}$
@@
$Pr\left[deg(v)\leq(1-\frac{1}{15})C_{p}\log n\right]\leq\exp\left(-\frac{C_{p}\log n}{675}\right)\leq\frac{1}{2n^{3}}$
Now, taking union bound over all vertices, we conclude that for every vertex $v$ it has $(1-\frac{1}{15})C_{p}\log n\leq deg(v)\leq(1+\frac{1}{15})C_{p}\log n$ with probability at least $1-\frac{1}{n^{2}}$ and thus with that probability $\frac{\Delta(G)}{\delta(G)}\leq\frac{1+\frac{1}{15}}{1-\frac{1}{15}}=1+\frac{1}{7}$
@@
Proof sketch..
The idea of the proof is to express $\overline{A+DE}$ as a sum of $\overline{A}$ and matrices with the small spectral norms and then apply a corollary of Weyl’s theorem, that is for $n\times n$ matrices $M_{1}$ and $M_{2}$
@@
The full proof can be found in Appendix C. ∎
Finally, using Lemmas 4 and 5 alongside the Theorem 3 we show that w.h.p. regularizing an Erdős–Rényi graph $G(n,p)$ by adding self-loops for every node to reach $\Delta(G)$ does not ruin its expansion properties.
Lemma 6.
@@
Proof.
Let $A$ be the adjacency matrix of $G$, $D$ be the degree matrix of $G$ and $E$ be the $n\times n$ diagonal matrix with entries $E_{ii}=\frac{\Delta(G)}{deg(v_{i})}-1$. By the Lemma 4, with probability $1-O(\frac{1}{n^{2}})$, $G$ has $\frac{\Delta}{\delta}\leq 1+\frac{1}{7}$ and therefore, $E_{ii}\leq\frac{1}{7}$ for all $i\in[n]$ with probability $1-O(\frac{1}{n^{2}})$.
Now, to each vertex $v$ in $G$, add $\Delta-deg(v)$ self-loops to obtain a graph $G^{\prime}$. Clearly, $G^{\prime}$ is $\Delta(G)$-regular. The adjacency matrix of $G^{\prime}$ will then be $A+DE$ and hence, applying Lemma 5 we deduce that $\lambda_{2}(G^{\prime})\leq\lambda_{2}(G)+\frac{6}{7}$.
By the Theorem 3, we know that with probability $1-O(\frac{1}{n^{2}})$, $\lambda_{2}(G)\leq\frac{C}{\sqrt{p(n-1)}}$ for some constant $C$, which for large enough $n$ is less than $\frac{1}{14}$, thus $\lambda_{2}(G^{\prime})\leq\frac{13}{14}$. ∎
4.3. Success of Multiple COBRAs
When multiple COBRA walks run in parallel, a given node might send multiple tokens along the single edge in one round. That is not feasible due to the congestion, hence we allocate $4\sqrt{\log n}$ rounds for each node to distribute its tokens. One phase corresponds to one round of individual COBRA.
Definition 0.
We say that a phase is $4\sqrt{\log n}$ rounds.
The quantity of $4\sqrt{\log n}$ is chosen so that every node manages to distribute its tokens w.h.p. Though, it is not guaranteed. To refer to this aspect of our algorithm, we introduce the following term.
Definition 0.
We say that multi-COBRA passes a phase if every node decides to send no more than $4\sqrt{\log n}$ tokens along a single edge in that phase.
We now formally state that the concept of a phase helps to avoid congestion and we justify the choice of the quantity $4\sqrt{\log n}$.
Lemma 9.
With probability $1-O(\frac{\log n}{n^{8}})$ multi-COBRA passes all $T$ phases.
Proof.
Let us fix a phase and a vertex $v$. By the design of the algorithm, $v$ has $t\leq 2\delta(G)$ tokens to send in the phase and $\Delta(G)$ adjacent edges. Let us fix an edge $e$ and define an indicator variable $\chi_{i}$ that is $1$ in case the $i$-th token is sent along $e$ and 00 otherwise. Denote $t_{e}=\sum\limits_{i\in[t]}\chi_{i}$. This way, the expected number of tokens sent along $e$ in one phase is $E\left[t_{e}\right]\leq t\cdot\frac{1}{\Delta(G)}\leq 2$. Now, by Chernoff’s bounds
$Pr\left[t_{e}\geq(1+4\sqrt{\log n})\cdot 2\right]\leq\exp\left(-\frac{16\log n\cdot 2}{3}\right)\leq\frac{1}{n^{10}}$
Since the degree of each node is at most $n$, there are at most $n^{2}$ edges. Hence, taking a union bound over all edges, we obtain that among every edge at most $4\sqrt{\log n}$ tokens are sent per phase with probability at least $1-\frac{1}{n^{8}}$. Finally, taking a union bound over all phases, we get that multi-COBRA passes all of them with probability $1-O(\frac{\log n}{n^{8}})$. ∎

### v2

@@ 2 Related Work
Despite these applications, to the best of our knowledge, the branching random walk remains underexplored in distributed computing, and this work seeks to showcase its untapped potential.
3 Algorithm Overview
In this section, we provide a high-level overview of our algorithm, which consists of two major parts. First, we embed a virtual Erdős–Rényi graph $G(n,p)$ atop the underlying (physical or overlay) network $H$, and then we solve the problem on $G$. We explain the embedding procedure in Section 3.1, and from that point on, we focus solely on solving multi-message broadcast on an Erdős–Rényi graph. In Section 3.2, we overview COBRA - the main building block of our algorithm for random graphs, and in Section 3.3 we describe the algorithm itself.
3.1 Embedding a Random Graph
In this section, we describe how to embed a random graph atop a given expander. A related technique was used by Ghaffari et al. [25], where it suffices to embed a sparse Erdős–Rényi graph with expected degree $O(\log n)$. In our setting, however, it is crucial to preserve the minimum degree of the host graph. To this end, we reuse Lemma 3.2 from [25], but we also introduce additional ideas of node groups and rejection sampling to meet our stronger requirements.
Embedding a graph $G$ atop a host graph $H$ involves creating virtual nodes $V(G)$ and establishing edges $E(G)$ between them so that
  • Every virtual node $u\in V(G)$ is simulated by some physical node $\mathrm{host}(u)\in V(H)$.
  • If a virtual node $u\in V(G)$ sends a message to $v\in V(G)$ along the edge $(u,v)\in E(G)$, this should be simulated by $\mathrm{host}(u)$ sending the same message to $\mathrm{host}(v)$ via some path in $H$.
Our construction will guarantee that each round of communication in $G$ can be simulated in $O(\tau_{mix}(H)\cdot\log n)$ rounds in $H$ and that $\delta(G)$ will be close to $\delta(H)$. This gives us the following “lifting”.
Theorem 3.1.
Consider a graph $H$ with $|V(H)|=n$. Assume there exists an algorithm that solves multi-message broadcast on an Erdős–Rényi graph $G(n^{\prime},\frac{\delta(H)}{n^{\prime}})$ in $R$ rounds w.h.p. for some $n^{\prime}=poly(n)$. Then there exists an algorithm that solves multi-message broadcast on $H$ in $O(R\cdot\tau_{mix}\cdot\log n+\tau_{mix}\cdot\log^{2}n)$ rounds w.h.p.
This round complexity consists of three terms: $O(R\cdot\tau_{mix}\cdot\log n)$ for simulating an algorithm, $O(\tau_{mix}\cdot\log^{2}n)$ rounds for constructing an embedding knowing $\tau_{mix}$, and $O(\tau_{mix}\cdot\log^{2}n)$ for estimating the $\tau_{mix}$. Below, we describe both the embedding and estimation parts.
Embedding Construction. We now outline how to construct the embedding of an Erdős–Rényi graph atop the expander $H$, assuming nodes have an estimate of $\tau_{mix}$.
We need to define a set of nodes of $G$. As an auxiliary concept, we start by defining sub-nodes. First, rethink $H$ as being bidirectional, that is, replace every undirected edge $(u,v)\in E(H)$ with two directed edges $(u,v)$ and $(v,u)$. Now, we associate a sub-node with each directed edge, and we say that a node $u\in V(H)$ simulates all the sub-nodes of its outgoing edges. Next, for each $u\in V(H)$, we arbitrarily group sub-nodes $u$ is simulating into groups of size exactly $\delta(H)$. This might leave some residual sub-nodes that will not be assigned to any group; we call those inactive, and others are called active. Note that at most half of all sub-nodes can be inactive. Finally, we define nodes in $G$ to be the aforementioned groups, and we make a node $u\in V(H)$ simulate a node $v\in V(G)$ if it simulates the sub-nodes in the respective group.
To establish edges in $G$, we launch lazy random walks of length $\tau_{mix}(H)$ from each active sub-node. The following lemma guarantees that those can be executed in parallel in only $O(\tau_{mix}(H)\cdot\log n)$ rounds w.h.p.
Lemma 3.2 (Lemma 2.5 [25]).
Let $H=(V,E)$ be an $n$-node graph. Suppose we wish to perform $T=n^{O(1)}$ steps of a collection of independent lazy random walks in parallel. If each node $v\in V$ initiates at most $d_{H}(v)$ walks, then w.h.p., the $T$ steps of all walks can be performed in $O(T\cdot\log n)$ rounds in a distributed setting.
If a random walk that started from a sub-node $a$, terminates at an active sub-node $b$, we call such a walk successful and we propagate the “success” message back to $a$ by simply executing the walk in reverse. This establishes the edge between $\mathrm{group}(a)\in V(G)$ and $\mathrm{group}(b)\in V(G)$. If, on the other hand, the walk from $a$ terminates at an inactive node, we call such a walk failed and we propagate the “failure” message back to $a$. Sub-nodes that received a “failure” message retry the same process again. The following lemma shows that only a few retries are needed.
Lemma 3.3.
After at most $O(\log n)$ retries, all sub-nodes will execute a successful random walk w.h.p.
Proof 3.4.
By the definition of a mixing time, a random walk from a given sub-node is distributed (almost) uniformly among all the sub-nodes. Hence, given that at least half of all sub-nodes are active, a probability of success in one round is at least $1/2$ (minus a negligible term due to the “almost” uniformity). Therefore, the probability of not succeeding once after $100\log n$ trials is no more than $\frac{1}{n^{10}}$. Applying the union bound over all sub-nodes completes the proof.
Once paths are established, communication along $(u,v)\in E(G)$ is simulated by routing a message along the corresponding walk path in $H$, which, by Lemma 3.2, can be done in $O(\tau_{mix}\cdot\log n)$ rounds.
We remark that the process described above is equivalent to sampling with replacement $\delta$ neighbors for each node $v\in V(G)$ independently (almost) uniformly at random. While this is not a canonical definition of an Erdős–Rényi graph, the resulting distribution is equivalent, modulo a negligible inverse-polynomially small probability.
On Distributed Estimation of Mixing Time. Note that in order to implement this simulation, nodes do not need prior knowledge of $\tau_{mix}$. Instead, they can estimate the mixing time of the network using the decentralized algorithm of Kempe and McSherry [36], which runs in $O(\log^{2}n\cdot\tau_{mix})$ rounds (without prior knowledge of $\tau_{mix}$). Their algorithm can be used to estimate the second principal eigenvalue $\lambda$ of the transition matrix, which relates to the mixing time via the following inequality:
Theorem 3.5 ([48]).
Given a graph on $n$ nodes with mixing time $\tau_{mix}$ and second principle eigenvalue $\lambda$, it holds that
$\left(\frac{1}{1-\lambda}-1\right)\log n\leq\tau_{mix}\leq 2\log n\cdot\frac{1}{1-\lambda}.$
Combining all together, we can prove Theorem 3.1.
Proof 3.6 (Proof of Theorem 3.1).
First, estimate the mixing time $\tau_{mix}(H)$ of the graph $H$ using the approach from [36]. This takes $O(\tau_{mix}(H)\cdot\log^{2}n)$ rounds.
Let the nodes of $H$ discover $\delta(H)$. This can be done through building a BFS tree in $H$ within $O(D(H))=O(\tau_{mix}(H))$ rounds.
Knowing $\tau_{mix}(H)$ and $\delta(H)$, compute an embedding as described above. Make the source node for a multi-message broadcast in $G$ any node in $G$ that is simulated by the source of an original problem instance in $H$. Solve the problem in $G$ in $R$ rounds, simulating each round of $G$ in $O(\tau_{mix}(H)\cdot\log n)$ rounds in $H$.
3.2 Coalescing-Branching Random Walk
@@
We now overview the main primitive for our algorithm for an Erdős–Rényi graph – the COalescing-BRAnching Random Walk (COBRA walk). COBRA walk was first introduced by Dutta et al. [15] in their work “Coalescing-Branching Random Walks on Graph” [15], with subsequent refinements presented in [12, 41, 6]. The COBRA walk is a generalization of the classical random walk, defined as follows: At round 00, a source node $s\in V$ possesses a token. At round $r$, each node possessing a token selects $\kappa$ of its neighbors uniformly at random, sends a token copy to each of them, and these neighbors are said to possess a token at round $r+1$. Here, $\kappa$, referred to as the branching factor, can be generalized to any positive real number [12]. When $\kappa=1$, the COBRA walk reduces to the classical random walk. From now on, we consider $\kappa$ to always be $2$. It is important to note that if a node receives multiple token copies in a round, it behaves as if it has received only one token; it will still choose $\kappa$ neighbors uniformly at random. This property, where received token copies coalesce at a node, gives the primitive its name.
@@
Cooper et al. [12] studied the cover time $T$ of the COBRA walk on regular expanders and obtained the following theorem which we use in our result
Theorem 3.7 (Cooper et al. [12]).
@@
Let $G$ be a connected $n$-vertex regular graph. Let $\lambda_{2}$ be the second largest eigenvalue (in the absolute value) of the normalized adjacency matrix of $G$. Then the following holds with probability at least $1-O(\frac{1}{n^{2}})$.
@@
We point out that in [12], the bound on probability is $1-O(\frac{1}{n})$, though the analysis, which is based on Chernoff bounds, can be adapted so that the probability is $1-O(n^{C})$ for any constant $C$ and the cover time is only multiplied by a constant.
In the upcoming analysis of our result, we will make sure that $\lambda_{2}$ is no more than $\frac{13}{14}$ w.h.p. Let $C_{T}$ be a sufficiently large constant so that a COBRA walk covers a regular graph with $\lambda_{2}\leq\frac{13}{14}$ with probability at least $1-O(\frac{1}{n^{2}})$ in $C_{T}\cdot\log n$ rounds. From now on, we define $T$ to be $C_{T}\log n$.
3.3 Random Graph Algorithm Description
The algorithm to solve the multi-message broadcast on an Erdős–Rényi graph $G(n,p)$ proceeds through the following steps:
  1. 1.
    Building a BFS Tree and Gathering Information: Nodes construct a BFS tree rooted at the source node $s$. Using the tree, every node learns the total number of nodes $|V|$, the minimum degree $\delta$, and the maximum degree $\Delta$. This step takes $O(D)$ rounds and does not require any prior knowledge of the graph topology.
  2. 2.
    Regularizing the Graph: Each node $v$ adds $\Delta-\textit{deg}(v)$ self-loops to its adjacency list to make the graph regular. In our analysis, we will show that each node adds a relatively small number of self-loops. This operation is purely local and requires no communication between nodes.
  3. 3.
    Constructing Spanning Subgraphs through Multiple COBRA Walks: The source node $s$ initiates $\delta$ COBRA walks by creating $\delta$ tokens labeled $1$ through $\delta$. When a token from the $i$-th COBRA walk is sent along an edge $e:\ (u,v)$, nodes $u$ and $v$ mark $e$ as part of the $i$-th subgraph. Note that a single edge may belong to multiple subgraphs.
    When running multiple COBRA walks simultaneously, congestion can occur if a node attempts to send multiple tokens from different COBRA walks along the same edge in a single round. Since only one token can traverse an edge per round, this creates a bottleneck that needs to be managed. To address this, we organize the process into phases, where each phase consists of $2$ rounds. In each phase, spanning rounds $\{2r,2r+1\}$, every node $u$ distributes two tokens for each COBRA walk whose token(s) it received during the previous phase. Since there are $\delta$ COBRA walks, node $u$ could have received at most $p\leq\delta$ distinct tokens. Let these tokens be denoted by $t_{1},\ldots,t_{p}$. Node $u$ then distributes each token twice, as follows: it enumerates its neighbors as $v_{1},\ldots,v_{\ell}$, with $\ell\geq\delta$, generates two independent random permutations $\sigma_{1},\sigma_{2}\in S_{\ell}$, and sends token $t_{i}$ to neighbor $v_{\sigma_{1}(i)}$ in round $2r$ and to neighbor $v_{\sigma_{2}(i)}$ in round $12r+$ 1. Consequently, for any fixed COBRA walk $i$ and any node $u$, the token belonging to the $i$-th COBRA walk is sent to a neighbor of $u$ chosen uniformly at random, independently of the other tokens in that same COBRA walk. (Different cobra walks, however, need not be independent.)
@@ 3.3 Random Graph Algorithm Description
  4. Parallel execution. For the rest of the algorithm, we run protocols on all the constructed subgraphs in parallel. To achieve this despite potential congestion (recall that each edge may belong to multiple subgraphs), we again organize the execution into phases. Now, each phase spans $2T$ rounds, ensuring that messages sent along any shared edge are distributed across the protocols without conflict. Specifically, if a protocol would send a message along an edge in a particular round when executed independently, all such messages from different subgraphs are scheduled within the same phase. This phased execution allows all protocols to proceed in parallel while respecting the edge capacity.
  5. 4.
    Constructing Tree Packings: The source $s$ initiates a BFS on each subgraph to transform it into a tree. By the end of this step, the algorithm constructs a tree packing $\{T_{i}\}_{i\in[\delta]}$. This step takes the number of phases that is at most the height of the highest tree, that is at most $O(T)$ phases, and hence $2T\cdot O(T)=O(T^{2})$ rounds.
  6. 5.
    Distributing Messages: The source node $s$ evenly divides the set of messages $M$ across the $\delta$ trees, ensuring that each tree receives $\frac{k}{\delta}$ messages. These messages are then downcasted along the trees one by one. To broadcast $\frac{k}{\delta}$ messages in each tree of height $O(T)$ in parallel, one needs $O(\frac{k}{\delta}+T)$ phases, that is $O(T\cdot(\frac{k}{\delta}+T))$ rounds.
4 Proof
In this section, we formally state our results and their auxiliaries for the Erdős–Rényi graph. We start by providing a high-level overview.
4.1 Proof Outline
The proof proceeds in three main steps. First, we argue that making a random graph regular by adding self-loops does not significantly affect its expansion properties. To this end, we rely on a result of Hoffman et al. [33], which shows that a random graph is a good expander with high probability. We also use standard Chernoff bound arguments to establish that random graphs are nearly regular, meaning the ratio between maximum and minimum degrees is close to one. Finally, we invoke a classical consequence of Weyl’s inequality, which ensures that small perturbations to the diagonal of a matrix only slightly affect its eigenvalues. Together, these ingredients imply that regularizing the graph preserves its spectral expansion up to a small error.
The second step is to prove that each individual COBRA walk successfully covers the entire network. The permutation trick allows us to claim that though COBRA walks are not independent, their marginal distributions stay as if they were. Using this, Theorem 3.7 by Cooper et al., together with a union bound, guarantees that all COBRA walks complete successfully within $O(\log n)$ phases w.h.p.
In the final step, we argue that the algorithm produces a tree packing with size $\delta(G)$, height $O(\log n)$, and weight $O(\log n)$. This tree packing allows for broadcasting all messages in $O(\log^{2}n+\log n\cdot\frac{k}{\delta(G)})$ rounds. This is optimal up to additive $\log^{2}n$ and multiplicative $\log n$, as $\frac{k}{\delta(G)}$ provides a natural lower bound for the broadcast time.
4.2 Introducing Regularity while Maintaining Expansion
@@
Theorem 4.3 (Hoffman et al. [33]).
For a positive constant $C$ and $p\geq\frac{C\log n}{n}$, consider an Erdős–Rényi graph $G(n,p)$. Then, with probability at least $1-O(\frac{1}{n^{C-1}})$, we have $\lambda_{2}(G)=O(\frac{1}{\sqrt{pn}})$.
The following Lemma shows that with high probability, an Erdős–Rényi graph $G(n,p)$ is almost regular.
Lemma 4.4.
@@
Proof 4.5.
Let us fix a vertex $v$ and consider the number of its incident edges. For each potential edge $e_{i}$, $i\in[n-1]$ let us introduce an indicator variable $\chi_{i}$ which is equal to $1$ if the edge exists and to 00 if it does not. The number of edges $v$ has is then $\sum\limits_{i\in[n-1]}\chi_{i}$. The expectation of that is $p(n-1)$, and applying Chernoff bounds, we get the following bounds on the degree of $v$
$Pr\left[\textit{deg}(v)\geq(1+\frac{1}{15})C_{p}\log n\right]\leq\exp\left(-\frac{C_{p}\log n}{675}\right)\leq\frac{1}{2n^{3}},$
@@
$Pr\left[\textit{deg}(v)\leq(1-\frac{1}{15})C_{p}\log n\right]\leq\exp\left(-\frac{C_{p}\log n}{675}\right)\leq\frac{1}{2n^{3}}.$
Now, taking union bound over all vertices, we conclude that for every vertex $v$ it holds that $(1-\frac{1}{15})C_{p}\log n\leq\textit{deg}(v)\leq(1+\frac{1}{15})C_{p}\log n$ with probability at least $1-\frac{1}{n^{2}}$. Thus, with probability $1-\frac{1}{n^{2}}$, we have that $\frac{\Delta(G)}{\delta(G)}\leq\frac{1+\frac{1}{15}}{1-\frac{1}{15}}=1+\frac{1}{7}$
@@
Proof 4.7 (Proof sketch.).
The idea of the proof is to express $\overline{A+DE}$ as a sum of $\overline{A}$ and matrices with the small spectral norms, and then apply a corollary of Weyl’s Theorem [54], that is for $n\times n$ matrices $M_{1}$ and $M_{2}$, it holds that
@@
The full proof can be found in Appendix D.
Finally, using Lemmas 4.4 and 4.6 alongside Theorem 4.3, we show that w.h.p., regularizing an Erdős–Rényi graph $G(n,p)$ by adding self-loops for every node to reach $\Delta(G)$ does not ruin its expansion properties.
Lemma 4.8.
@@
Proof 4.9.
Let $A$ be the adjacency matrix of $G$, $D$ be the degree matrix of $G$ and $E$ be the $n\times n$ diagonal matrix with entries $E_{ii}=\frac{\Delta(G)}{\textit{deg}(v_{i})}-1$. By the Lemma 4.4, with probability $1-O(\frac{1}{n^{2}})$, $G$ has $\frac{\Delta}{\delta}\leq 1+\frac{1}{7}$ and therefore, $E_{ii}\leq\frac{1}{7}$ for all $i\in[n]$ with probability $1-O(\frac{1}{n^{2}})$.
Now, to each vertex $v$ in $G$, add $\Delta-\textit{deg}(v)$ self-loops to obtain a graph $G^{\prime}$. Clearly, $G^{\prime}$ is $\Delta(G)$-regular. The adjacency matrix of $G^{\prime}$ will then be $A+DE$ and hence, applying Lemma 4.6 we deduce that $\lambda_{2}(G^{\prime})\leq\lambda_{2}(G)+\frac{6}{7}$.
By the Theorem 4.3, we know that with probability $1-O(\frac{1}{n^{2}})$, $\lambda_{2}(G)\leq\frac{C}{\sqrt{p(n-1)}}$ for some constant $C$, which for large enough $n$ is less than $\frac{1}{14}$, thus $\lambda_{2}(G^{\prime})\leq\frac{13}{14}$.
4.3 Success of Multiple COBRAs
Let us recall the permutation trick used to avoid congestion among multiple COBRA walks. To be able to run multiple COBRA walks in parallel, we partition the execution into phases, each lasting $2$ rounds. In each phase, spanning rounds ${2r,2r+1}$, every node $u$ distributes two tokens for each COBRA walk whose token(s) it received during the previous phase. Since there are $\delta$ COBRA walks, node $u$ could have received at most $p\leq\delta$ distinct tokens. Let these tokens be denoted by $t_{1},\ldots,t_{p}$. Node $u$ then distributes each token twice, as follows: it enumerates its neighbors as $v_{1},\ldots,v_{\ell}$, with $\ell\geq\delta$, generates two independent random permutations $\sigma_{1},\sigma_{2}\in S_{\ell}$, and sends token $t_{i}$ to neighbor $v_{\sigma_{1}(i)}$ in round $2r$ and to neighbor $v_{\sigma_{2}(i)}$ in round $12r+$ 1.
The claim below summarizes the properties resulting from the procedure described:
Claim 1.
For each COBRA walk, a phase corresponds exactly to a round in the execution where this COBRA walk runs in isolation. In particular, for a given COBRA walk, every token is sent independently of other tokens of this walk and to a uniformly chosen neighbor.